OpenStudy (anonymous):

what is the absolute value of (x-2)/(x-2) limit 2

OpenStudy (anonymous):

\[\frac{x-2}{x-2}=1\] for all \(x\) except \(x=2\) so the limit is \(1\)

OpenStudy (anonymous):

oooh i bet it is this \[\frac{|x-2|}{x-2}\] am i right?

OpenStudy (anonymous):

no the absolute sign is at the bottom

OpenStudy (anonymous):

at the bottom but not at the top?

OpenStudy (anonymous):

yes that's correct

OpenStudy (anonymous):

ok then \[\frac{x-2}{|x-2|}\] is really only two numbers if \(x>2\) it is \(\frac{x-2}{x-2}=1\) but if \(x<2\) you have \(\frac{x-2}{|x-2|}=\frac{x-2}{-x+2}=-1\)

OpenStudy (anonymous):

in other words, it is a fancy way of writing \[f(x) = \left\{\begin{array}{rcc} -1 & \text{if} & x <2\\ 1& \text{if} & x >2 \end{array} \right. \]

OpenStudy (anonymous):

since evidently \(-1\neq 1\) this limit is undefined 

OpenStudy (anonymous):

can you please show me the workings, what I am seeing is only the answer

OpenStudy (anonymous):

can you not see what i wrote?

OpenStudy (anonymous):

if you refresh your browser you should be able to see that math that i wrote if you cannot see it now

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw